Let a and β be multi-indices such that β ≤ a. Prove that ∂^(β) x^(a) = (a,β) β!x^(a - β) if β≤a
1
Expert's answer
2019-10-04T10:05:38-0400

First we prove statement for one-dimensional case, that is "(x^a)^{(b)}=\binom{a}bb!x^{a-b}", where "b\le a".

We apply mathematical induction on "b\le a".

Indeed, "(x^a)^{(0)}=\binom{a}00!x^{a-0}=x^a". Suppose that "(x^a)^{(b)}=\binom{a}bb!x^{a-b}" for all "b\le k" , where "0\le k\le a".

If "k=a" , then our statement is proved. If "k<a" , then "(x^a)^{(k+1)}=\bigl((x^a)^{(k)}\bigr)'" .

By the induction hypothesis we have "(x^a)^{(k)}=\binom{a}kk!x^{a-k}" , so "(x^a)^{(k+1)}=\bigl((x^a)^{(k)}\bigr)'=\bigl(\binom{a}kk!x^{a-k}\bigr)'="

"=\binom{a}kk!\cdot(a-k)x^{a-k-1}"

Consider "\binom{a}kk!\cdot(a-k)" .

"\binom{a}kk!\cdot(a-k)=\frac{a!}{k!(a-k)!}k!(a-k)=\frac{a!}{(a-k-1)!}="

"=\frac{a!}{(a-k-1)!(k+1)!}(k+1)!=\binom{a}{k+1}(k+1)!"

So "(x^a)^{(k+1)}=\binom{a}kk!\cdot(a-k)x^{a-k-1}=\binom{a}{k+1}(k+1)!x^{a-k-1}" .

By the mathematical induction we obtain "(x^a)^{(b)}=\binom{a}bb!x^{a-b}" for all "b\le a" .

Consider multi-indices. We prove by induction on "n" that "\frac{\partial^{|\beta|} x^\alpha}{\partial x^\beta}=\binom{\alpha}{\beta}\beta!x^{\alpha-\beta}" , where "x=(x_1,\ldots,x_n)", "\alpha=(\alpha_1,\ldots,\alpha_n)", "\beta=(\beta_1,\ldots,\beta_n)" and "\beta\le\alpha".

We proved above this statement for "n=1". Suppose that the statement is true for "n=k", where "k\ge 1" .

Let "x=(x_1,\ldots,x_k,x_{k+1})", "\alpha=(\alpha_1,\ldots,\alpha_k,\alpha_{k+1})", "\beta=(\beta_1,\ldots,\beta_k,\beta_{k+1})" and "\beta\le\alpha".

Introduce new denotations: "x_0=(x_1,\ldots,x_k)", "\alpha_0=(\alpha_1,\ldots,\alpha_k)", "\beta_0=(\beta_1,\ldots,\beta_k)".

We have "\frac{\partial^{|\beta|} x^\alpha}{\partial x^\beta}=\frac{\partial^{|\beta_0|+\beta_{k+1}} x_0^{\alpha_0}x_{k+1}^{\alpha_{k+1}}}{\partial x_0^{\beta_0}\partial x_{k+1}^{\beta_{k+1}}}=\frac{\partial^{\beta_{k+1}}}{\partial x_{k+1}^{\beta_{k+1}}}\bigl(\frac{\partial^{|\beta_0|} x_0^{\alpha_0}x_{k+1}^{\alpha_{k+1}}}{\partial x_0^{\beta_0}}\bigr)=" "=\frac{\partial^{\beta_{k+1}}}{\partial x_{k+1}^{\beta_{k+1}}}\bigl(x_{k+1}^{\alpha_{k+1}}\frac{\partial^{|\beta_0|} x_0^{\alpha_0}}{\partial x_0^{\beta_0}}\bigr)=\frac{\partial^{|\beta_0|} x_0^{\alpha_0}}{\partial x_0^{\beta_0}}\frac{\partial^{\beta_{k+1}} x_{k+1}^{\alpha_{k+1}}}{\partial x_{k+1}^{\beta_{k+1}}}"Since "\beta\le\alpha", we obtain "\beta_0\le\alpha_0" and "\beta_{k+1}\le\alpha_{k+1}", so "\frac{\partial^{\beta_{k+1}} x_{k+1}^{\alpha_{k+1}}}{\partial x_{k+1}^{\beta_{k+1}}}=\binom{\alpha_{k+1}}{\beta_{k+1}}\beta_{k+1}!x_{k+1}^{\alpha_{k+1}-\beta_{k+1}}" and by induction hypothesis we have "\frac{\partial^{|\beta_0|} x_0^{\alpha_0}}{\partial x_0^{\beta_0}}=\binom{\alpha_0}{\beta_0}\beta_0!x_0^{\alpha_0-\beta_0}" . We obtain "\frac{\partial^{|\beta|} x^\alpha}{\partial x^\beta}=\frac{\partial^{|\beta_0|} x_0^{\alpha_0}}{\partial x_0^{\beta_0}}\frac{\partial^{\beta_{k+1}} x_{k+1}^{\alpha_{k+1}}}{\partial x_{k+1}^{\beta_{k+1}}}=" "=\binom{\alpha_0}{\beta_0}\beta_0!x_0^{\alpha_0-\beta_0}\cdot\binom{\alpha_{k+1}}{\beta_{k+1}}\beta_{k+1}!x_{k+1}^{\alpha_{k+1}-\beta_{k+1}}=" "=\binom{\alpha_0}{\beta_0}\binom{\alpha_{k+1}}{\beta_{k+1}}\beta_0!\beta_{k+1}!x_0^{\alpha_0-\beta_0}x_{k+1}^{\alpha_{k+1}-\beta_{k+1}}=\binom{\alpha}{\beta}\beta!x^{\alpha-\beta}" By the mathematical induction we obtain "\frac{\partial^{|\beta_0|} x_0^{\alpha_0}}{\partial x_0^{\beta_0}}=\binom{\alpha}{\beta}\beta!x^{\alpha-\beta}" for all "n", "x=(x_1,\ldots,x_n)", "\alpha=(\alpha_1,\ldots,\alpha_n)", "\beta=(\beta_1,\ldots,\beta_n)", where "\beta\le\alpha" .